# Learning to Solve Multi-Objective Routing Problems on Multigraphs

### v1

@@ Abstract
Learning-based methods for routing have gained significant attention in recent years, both in single-objective and multi-objective contexts. However, the multigraph setting, where multiple paths with distinct attributes can exist between destinations, has largely been overlooked, despite its high practical relevancy. In this paper, we introduce two neural approaches to address multi-objective routing on multigraphs. Our first approach works directly on the multigraph, by autoregressively selecting edges until a tour is completed. On the other hand, our second model first prunes the multigraph into a simple graph and then builds routes. We validate both models experimentally and find that they demonstrate strong performance across a variety of problems, including the Traveling Salesman Problem (TSP) and Capacitated Vehicle Routing Problem (CVRP)111Our code will be made publicly available upon acceptance of the paper..
@@ 1 Introduction
Routing problems, including the well-known Traveling Salesman Problem (TSP), are central for optimizing operations in logistics, delivery and beyond. However, these problems are generally NP-hard and difficult to solve exactly in practical, real-time settings. While exact solvers provide high-quality solutions, they do not scale well. At the same time, hand-crafted heuristics are often efficient but time-consuming to design and typically lack generalizability across problem variants.
Recently, neural solvers have emerged as a promising alternative, offering speed, accuracy, and broad applicability across routing problems (Kool, van Hoof, and Welling 2018; Ye et al. 2024; Drakulic, Michel, and Andreoli 2025). However, most learning-based approaches focus on Euclidean settings, where customers are points in a plane and travel costs are based on the Euclidean distance. Real-world road networks are far more complex and travel cost is usually asymmetric, affected by elevation, one-way streets and traffic. Further, when there are multiple conflicting criteria of importance, there can be multiple paths between customers that cannot be chosen between apriori.
In such cases, a directed multigraph, allowing multiple, distinct paths between the same locations, is a more realistic and powerful input-structure. In fact, using this representation has empirically been shown to increase solution quality significantly compared to the simple graph representation (Ben Ticha et al. 2017). Yet this comes at the cost of increased complexity, underscoring the need for new methods capable of handling multigraph-based routing efficiently.
The challenge intensifies in Multi-Objective (MO) routing problems, such as the Multi-Objective TSP (MOTSP), where several criteria (e.g., distance and time) must be optimized simultaneously. Even without multigraphs, MO problems are hard: not only because their single-objective counter-part is often NP-hard, but also because the number of Pareto optimal solutions can increase exponentially with the size of the problem (Ehrgott and Gandibleux 2000).
Pareto Set Learning (PSL) has recently emerged as a powerful framework for MO routing. By training offline on similar instances, PSL approximates the Pareto set quickly during inference. Moreover, since the framework is based on reinforcement learning, there is no need for oracle solvers for supervised training. These methods have shown strong results in Euclidean settings (Lin, Yang, and Zhang 2022; Fan et al. 2025), but there is currently no general PSL framework for routing on multigraphs or in non-Euclidean settings, despite their higher practical relevance.
We attempt to bridge this gap. Concretely, our main contributions are:
@@
We extend PSL to handle multigraph inputs, where the model selects not only the next node but also the path taken. To our knowledge, this is the first neural solver designed for such input structures. Although we treat multi-objective optimization, we hypothesize that our approach can be useful in single-objective constrained and stochastic cases too.
As a byproduct, our model constitutes the first general-purpose, multi-objective solver capable of handling a variety of routing problems on asymmetric graphs.
@@
In order to do this, we propose two separate models: One edge-based that is simple and accurate but relatively slow and one dual-task model that first sparsifies the multigraph into a simple graph and then selects routes. The latter is faster but more complex.
@@
Experimentally, we show that our approach achieves competitive results for several MO routing problems on asymmetric graphs and multigraphs, including the MOTSP, Multi-Objective Capacitated Vehicle Routing Problem (MOCVRP) and MOTSP with time-windows.
@@ 2.1 Learning for Routing
Motivated by the success of end-to-end learning rather than hand-crafted heuristics in fields such as image classification and language modeling, neural heuristics for routing have gained increasing attention during the last years. Broadly, methods can be classified by how they obtain the tours. Two prominent alternatives are autoregressive construction, based on iterative building of tours one node at a time, and non-autoregressive construction, that outputs heatmaps in a one-shot fashion for down-stream decoding with some search method.
Early works in the first category include Vinyals, Fortunato, and Jaitly (2015), which utilized an oracle TSP solver for supervised learning, and Bello et al. (2017), which introduced reinforcement learning for training. By using attention models and inherent symmetries in the problem, first Kool, van Hoof, and Welling (2018) and then Kwon et al. (2020) obtained close to 0 optimality gap on various routing problems. Regarding asymmetric problems, Kwon et al. (2021) proposed MatNet, a mixed-attention architecture working on bi-partite graphs. More recent works are Jin et al. (2023); Drakulic et al. (2023); Chen et al. (2024); Bi et al. (2024). These papers have improved results, can handle constraints better and can scale to even larger problems than previously. Another notable example is Drakulic, Michel, and Andreoli (2025), where one model learns to solve a multitude of combinatorial optimization problems, including examples from routing.
In the second category, some notable works are Joshi, Laurent, and Bresson (2019); Fu, Qiu, and Zha (2021); Qiu, Sun, and Yang (2022); Sun and Yang (2023). Typically, these methods perform worse than autoregressive approaches, but scale better to large instances due to lower space and time complexity. An interesting recent work combining the two approaches to solve large instances (up to nodes for the TSP) is Ye et al. (2024).
2.2 Multi-Objective Learning for Routing
The literature on multi-objective routing is more sparse than in the single-objective case. Some early works (Li, Zhang, and Wang 2021; Wu, Wang, and Zhang 2020) utilize multiple models, where one model is responsible for one preference between the objectives. This quickly becomes infeasible as the number of models required grows exponentially with the number of objectives (Ehrgott 2005). Pareto-set learning for multi-objective routing using a single model was proposed in Lin, Yang, and Zhang (2022). They used the hyper-networks of Navon et al. (2021) to condition the decoder on the preference. Recent papers that feature improved results with a similar conditioning method include Fan et al. (2025); Lu et al. (2024). An alternative approach is to use meta-learning to train a single model that can be quickly fine-tuned over a few steps to solve individual subproblems. Two representative works in this direction are Zhang et al. (2021); Chen et al. (2023a).
Regarding asymmetric problems in the context of multi-objective optimization, the only existing learning-based methods we are aware of are those proposed in Santiyuda et al. (2024) and Zhou, Kong, and Wang (2025). Notably, both of these methods are tailored to particular problem types and evaluated within narrowly defined domains. In contrast, our approach is designed with a very general framework, capable of handling a broad class of routing problems.
@@ 2.3 Routing on Multigraphs
In the multi-objective case, a multigraph is often a more suitable representation of a road network than a simple graph. In fact, this network model has become more and more common for MO shortest path problems. One solution is genetic algorithms, see for instance Beke, Weiszer, and Chen (2021) for a comprehensive benchmarking study. Other studies have proposed A∗-related algorithms, e.g., AMOA∗ (Weiszer, Burke, and Chen 2020) and MOMGA∗ (Liu, Chen, and Weiszer 2022), to efficiently obtain a Pareto set of shortest paths.
The multigraph framework has also been applied to routing problems, but to a lower extent. The first study in this regard was Garaix et al. (2010), who designed a simple heuristic and branch-and-price method for a single-objective on-demand transportation problem. Other works proposing either exact solvers or heuristics for single-objective constrained problems or multi-objective problems are Lai, Caliskan Demirag, and Leung (2016); Ben Ticha et al. (2019); Tikani, Setak, and Demir (2021). In addition, a particularly interesting study is Ben Ticha et al. (2017), where it is clearly concluded that a multigraph setup yields significantly lower costs compared to a simple graph representation for the capacitated vehicle routing problem with time-windows. On real-world instances, a solution cost reduction of up to is reported. At the same time, the computation time increases significantly under this perspective, and the authors highlight the need for new heuristics.
We believe learning-based methods can fill this niche, as they have proven to be efficient, effective and broadly applicable. To the best of our knowledge, this work presents the first learning-based routing approach to address problems on multigraphs.
3 Multi-Objective Optimization
A general multi-objective optimization problem can be written as
where and . Unlike the Single-Objective (SO) case, the solution to such a problem is not one objective value with corresponding optimal solutions but a set of values known as the Pareto Front (PF) with corresponding solutions in the Pareto Set (PS). Formally, we define the PS as the set
@@ 3 Multi-Objective Optimization
To obtain this set, a typical method is decomposition through scalarization. The idea is to translate the MO problem into several SO problems that can be solved in parallel or consecutively. Let such that and be a vector that controls the preference among the objectives. Then, the most straight-forward scalarization is the Linear Scalarization (LS) . The LS subproblem is then , which is a standard SO problem. Unfortunately, it is well-known that there may be solutions in the Pareto Set that do not solve an LS subproblem for any preference (Ehrgott 2005). The Chebyshev Scalarization has more favorable properties. Let be an ideal value for objective . Then the Chebyshev scalarized cost is
Given a solution in the PS, there always exists a preference and corresponding subproblem for which the solution is optimal (Choo and Atkins 1983).
In our context of multigraphs, it is interesting to consider how the problem simplifies under scalarization. We note that under LS, the MOTSP on a multigraph can be simplified to a simple graph by pruning edges that have sub-optimal LS cost compared to a parallel edge. Under Chebyshev scalarization or for more complicated problems this simple method will not work. Instead, we propose to learn which parallel edge is optimal while simultaneously learning to construct the optimal route.
@@ 4 GNN-Based Multigraph Solver
In the following, we detail two approaches based on Graph Neural Networks (GNNs) for solving MO multigraph routing problems. We call our method GNN-based Multigraph Solver (GMS) and propose an edge-based variant, which is slow but simple, as well as an architecture with two decoders, which is faster but more complex. Both of these utilize the Graph Edge Attention Network (GREAT) of Lischka et al. (2024) as the encoder, the reason being that it 1) Can handle input data structured as multigraphs; 2) Outputs edge-embeddings; 3) Has been applied in an autoregressive-framework and shown competitive performance for non-Euclidean SO problems. However, note that it is possible to use other architectures in the encoder, such as e.g., the anisotropic GNN of Sun and Yang (2023).
@@ 4.1 Edge-Based GMS
We visualize Edge-Based GMS (GMS-EB) in Figure 1. The encoder, consisting of GREAT-layers, outputs edge embeddings. Using these, the decoder constructs valid tours auto-regressively. As in Lin, Yang, and Zhang (2022); Navon et al. (2021) we utilize a hyper-network to generate the weights of the decoder conditioned on the current preference. The decoder is an edge-based variant of the Multi-Pointer (MP) decoder from Jin et al. (2023), which we detail in Appendix A. Given the context of the current node in the tour, the decoder outputs probabilities of traversing edges starting in that node.
Note that the underlying graphs in routing problems are usually complete. Hence there are edges in a graph of nodes. Moreover, the decoder needs to be run times to construct a full tour and we usually roll out sample trajectories. Hence an edge-based decoder scales according to compared to a node-based that scales according to . We find that this is a bottleneck in terms of memory requirements and inference time. Hence we propose a second architecture that utilizes a node-based decoder.
@@ 4.2 Dual Head GMS
Using GREAT, we can obtain node-embeddings by not applying (2) in the final layer. The problem is that the action of selecting a node is not sufficient, as a path selection is also needed. Thus, we need to ensure that the action of choosing a node uniquely defines which edge to choose, which is achieved by pruning parallel edges leaving a unique one between each node-pair. In general, choosing this edge is not trivial. Thus, we propose to learn the choice using a dual decoder setup.
We call this approach Dual Head GMS (GMS-DH) and it is visualized in Figure 2. The key is that we insert a second decoder before the final encoder layer. This decoder uses the intermediate edge-embeddings to select which parallel edge to keep. The selection is fed back into the encoder, which only propagates the embedding of the remaining edges and converts them to node-embeddings. We also find it beneficial to add a small number of standard transformer layers after the final GREAT-layer in order to obtain more expressive node-embeddings. We detail the architecture for the selection-decoder in Appendix A. For the routing-decoder, we use a preference-conditioned version of the Multi-pointer decoder (Jin et al. 2023). When applying GMS-DH to simple graphs, we simply remove the first decoder. Also note that this architecture ensures that the first layers only need to be run once for each instance, which reduces the run-time compared to selecting edges before encoding.
From a broader perspective, our approach in GMS-DH can be viewed as a mixture of non-autoregressive and autoregressive learning. As a drawback, the edge-choice can no longer depend on the routing context, instead it must be chosen before the roll-out is started.
@@ 4.3 Training Algorithm
Our first proposed model, GMS-EB, is trained using the Multi-Objective REINFORCE from Lin, Yang, and Zhang (2022). In short, given a preference and instance , the model parameterizes a policy . A tour is associated with a reward according to (1)
The objective is to maximize
Additionally, we form a baseline by rolling out tours for the same instance and averaging the reward, according to the POMO framework (Kwon et al. 2020), to stabilize the training. The policy gradient is utilized to update the model weights and it is approximated, given one preference , according to
@@
For the second model, the training is more complicated. We can view each head as an agent. The selection-head takes actions corresponding to which edge to choose between each node pair. It parameterizes a policy . The routing-head uses the selection, which we denote by . It parameterizes a policy and obtains a scalar reward according to (3). With a fixed , the objective for this head is to maximize
On the other hand, the task of the selection-head is to ensure that the edge selection is optimal. Let denote the optimal tour given edge selection . Then, define
@@
For easy reference, we outline the training of GMS-DH in Algorithm 1.
@@
In addition, for both models we utilize a curriculum learning approach to speed up training. We start by training on small graph sizes and then gradually increase the problem size. For GMS-DH in the multigraph setting, we also start by training the routing-head on simple graphs to ensure that the approximation (4) is accurate even during the initial epochs of multigraph training. Further details regarding the curriculum learning can be found in Appendix B.

### v2

# Neural Approaches for Multi-Objective Routing on Multigraphs

@@ Abstract
Learning-based methods for routing have gained significant attention in recent years, both in single-objective and multi-objective contexts. Yet, existing methods are unsuitable for routing on multigraphs, which feature multiple edges with distinct attributes between node pairs, despite their strong relevance in real-world scenarios. In this paper, we propose two graph neural network-based methods to address multi-objective routing on multigraphs. Our first approach operates directly on the multigraph by autoregressively selecting edges until a tour is completed. The second model first simplifies the multigraph via a learned pruning strategy and then performs routing on the resulting simple graph. We evaluate both models empirically and demonstrate their strong performance across a range of problems and distributions111Our code will be released publicly upon paper acceptance..
@@ 1 Introduction
Routing problems, including the well-known Traveling Salesman Problem (TSP), are central for optimizing operations in modern transport systems. A key assumption is often that an optimal path can be selected beforehand between each pair of locations. However, in reality, several paths may exist that cannot be chosen between a priori. This naturally leads to a multigraph representation. Notably, by planning on a multigraph rather than a simple graph, Ben Ticha et al. (2017) demonstrated a cost reduction of up to for the vehicle routing problem with time-windows. Similarly Lai, Caliskan Demirag, and Leung (2016) demonstrated average savings of circa for a heterogeneous vehicle routing problem. Such results can be pivotal in for instance last-mile-delivery, which accounts for of total supply chain cost according to Capgemini Research Institute (2023).
Yet, routing problems on simple graphs are already difficult to solve, especially in real-time, as they are generally NP-hard. The multigraph setup further increases this complexity. Moreover, such a framework is particularly interesting in the Multi-Objective (MO) case, where each edge may be associated with several conflicting cost criteria, such as time, distance or energy. MO routing is typically harder, as the number of Pareto optimal solutions can grow exponentially with the size of the problem (Ehrgott and Gandibleux 2000). Clearly, MO routing on multigraphs poses significant challenges, while remaining highly relevant for real-world applications.
Recently, however, learning-based methods have emerged as promising options for routing, offering speed, accuracy, and broad applicability across problems (Kool, van Hoof, and Welling 2018; Drakulic, Michel, and Andreoli 2025). Yet, current approaches are designed for simple graphs and not suited for the multigraph representation. This is especially true in the MO setting, where neural approaches have demonstrated excellent results on Euclidean instances (Lin, Yang, and Zhang 2022; Fan et al. 2025), but lack support even for asymmetric graphs, in spite of their higher practical relevance (Boyacı, Dang, and Letchford 2021).
We attempt to bridge the gap between learning-based methods for MO routing and accurate network representations. Concretely, our main contributions are:
@@
We design models to handle multigraph inputs, where not only the next node but also the edge must be selected. To our knowledge, we present the first neural solvers designed for such input structures. Although we deal with MO optimization, our approach can be highly useful in single-objective cases too.
Our models are also capable of handling a variety of MO routing problems on asymmetric graphs. In contrast, current neural MO methods are either problem-specific or designed exclusively for the Euclidean setting, where edge costs are symmetric and determined only by the node coordinates.
@@
We propose two separate models based on Graph Neural Networks (GNNs): One edge-based working directly on the multigraph and one dual-task model that first prunes the multigraph into a simple graph and then selects routes. The former is simple but slow while the latter is more complex but faster.
@@
Experimentally, we show that our approach achieves competitive results for several MO routing problems on asymmetric graphs and multigraphs.
@@ 2.1 Learning for Routing
Motivated by the success of end-to-end learning in fields such as image classification and language modeling, neural solutions for routing have gained increasing attention during the last years. Broadly, methods can be classified by how they obtain tours. Two prominent alternatives are autoregressive construction, based on iterative building of tours one node at a time, and non-autoregressive construction, that outputs heatmaps in a one-shot fashion for down-stream decoding with some search method. Examples from the first category include Vinyals, Fortunato, and Jaitly (2015); Kool, van Hoof, and Welling (2018); Kwon et al. (2020). Examples from the second category are Joshi, Laurent, and Bresson (2019); Fu, Qiu, and Zha (2021); Sun and Yang (2023). For a more comprehensive discussion of these methods, we refer the reader to Appendix A, which presents additional related work.
The literature on MO routing is more sparse than in the single-objective case. A key contribution was made by Lin, Yang, and Zhang (2022), who proposed an autoregressive construction approach using a single neural model to learn the Pareto set of solutions. Regarding asymmetric problems, the only existing learning-based methods we are aware of are those introduced by Santiyuda et al. (2024) and Zhou, Kong, and Wang (2025). However, both these approaches are tailored to specific problems and have been evaluated within narrowly defined domains. In contrast, our method is capable of addressing a broad class of routing problems and is evaluated across various problems and settings.
@@ 2.2 Routing on Multigraphs
The multigraph representation has become more and more common for MO shortest path problems during recent years (Weiszer, Burke, and Chen 2020; Beke, Weiszer, and Chen 2021). Such a framework has also been applied to combinatorial routing problems like Vehicle Routing Problems (VRPs), but to a lower extent. Garaix et al. (2010) performed the first study in this regard and later works include Lai, Caliskan Demirag, and Leung (2016); Ben Ticha et al. (2017); Tikani, Setak, and Demir (2021).
A common theme across these studies is that the multigraph representation leads to considerably improved solutions, while simultaneously increasing the complexity of these already difficult problems. Consequently, learning-based methods should be able to fill a clear need for new solvers in this area, as they have proven to be efficient, effective and broadly applicable. To the best of our knowledge, this work presents the first learning-based approach to address combinatorial routing problems such as the TSP and VRPs on multigraphs.
3 Problem Formulation
We define a general multi-objective routing problem as
where with . Here, represents a feasible route in the multigraph , encoded as a sequence of edges, and is the associated cost vector. Due to the combinatorial nature of the feasible solution space and the multiple conflicting objectives, the problem is generally NP-hard.
The solution to an MO routing problem is a set of objective values known as the Pareto Front (PF) with associated routes in the Pareto Set (PS). Formally, we define the PS as
@@ 3 Problem Formulation
To obtain this set, we utilize a typical method: decomposition through scalarization. The idea is to translate the MO problem into several subproblems that can be solved separately. Let such that and be a vector that controls the preference among the objectives. Then, under Linear Scalarization (LS), the single-objective cost is and the corresponding subproblem is . Unfortunately, it is well-known that there may be solutions in the PS that do not solve an LS subproblem for any preference (Ehrgott 2005). The Chebyshev scalarization has more favorable properties. Let be an ideal value for objective . Then, the Chebyshev scalarized cost is
Given a solution in the PS, there always exists a preference and corresponding Chebyshev subproblem for which the solution is optimal (Choo and Atkins 1983).
In our context of multigraphs, it is interesting to consider how problems simplify under scalarization. We highlight an important result for the Multi-Objective TSP (MOTSP), in which the total cost of a tour is the sum of the vectorial costs of the traversed edges. The proof of the following proposition is omitted due to its simplicity. Given a linear scalarization defined by a preference vector , define the reduced graph by removing any edge whose scalarized cost is strictly worse than that of a parallel edge.
Proposition 1.
Let denote the linearly scalarized cost and let be the set of feasible tours on the pruned graph . Then, the optimal value of the scalarized subproblem is preserved:
This illustrates that a multigraph representation is sometimes easy to handle using pruning. However, under Chebyshev scalarization or for more complex problems (e.g., with edge attributes that contribute nonlinearly to the cost), similar results might not be available. As such, we propose to learn which parallel edge is optimal while simultaneously learning to construct the optimal route for each preference . Moreover, by explicitly taking into account the multigraph representation when designing our models, we obtain better results and faster inference than if we pre-prune the multigraphs, even for the MOTSP.
@@ 4 GNN-Based Multigraph Solver
In the following, we propose two approaches based on GNNs for solving MO multigraph routing problems. We call our method GNN-based Multigraph Solver (GMS) and propose an edge-based variant constructing routes autoregressively directly on the multigraph as well as a variant with two decoders that first prunes and then constructs routes.
Both variants require an encoder that i) can handle input data structured as multigraphs and ii) works with and outputs edge embeddings. While there are multiple GNN architectures that fit this description, we utilize the Graph Edge Attention Network (GREAT) of Lischka et al. (2025), since it has been applied in an autoregressive framework and shown competitive performance for non-Euclidean single-objective routing problems.
@@ 4.1 Edge-Based GMS
Vanilla node-based construction is insufficient for our purpose, as it relies on predicting the next node in the route, while we also require an edge-selection between the current node and the next. Thus, we propose an end-to-end edge-based model that instead predicts the next edge, and thereby implicitly the next node. We call this Edge-based GMS (GMS-EB).
We visualize GMS-EB in Figure 1. The encoder, consisting of GREAT-layers, outputs edge embeddings. Using them, the decoder constructs valid tours autoregressively. Given the instance and incomplete route in construction step , the decoder selects edge with probability . Thus the probability of the whole route is
Here, is the weights of the model given the current preference . As Navon et al. (2021) and Lin, Yang, and Zhang (2022), we utilize a Multi-Layer Perceptron (MLP) hyper-network to preference-condition the decoder, while keeping the encoder preference-agnostic. Consequently we form the decoder weights according to and set
Regarding the decoder architecture, it is an edge-based variant of the Multi-Pointer (MP) decoder from Jin et al. (2023). We detail this component further in Appendix B.
Note that a multigraph with nodes and edges between each pair of nodes has edges in total. Moreover, the decoder must be run times to construct a full tour and we roll out sample trajectories according to the POMO framework (Kwon et al. 2020). Hence, an edge-based decoder scales as compared to a node-based one that scales as . Consequently, we propose a second architecture that employs a node-based decoder and thereby scales better to higher node counts in terms of memory requirements and runtime.
@@ 4.2 Dual Head GMS
For a node-based decoder, we need to ensure that the action of choosing a node uniquely defines which edge to choose. This is achieved by pruning parallel edges, leaving a unique one between each node pair. In general, choosing this edge is not trivial. As such, we propose to learn the choice using a dual decoder setup.
We call this approach Dual Head GMS (GMS-DH) and it is visualized in Figure 2. The key is that we insert another decoder before the final GREAT layer, which uses intermediate edge embeddings to select the parallel edge to keep. Let be the set of edges between the node pair . The selection-decoder produces the probability of retaining . As we treat each node pair independently and select one edge per pair, the probability of the joint selection of active edges is
Again, the weights are generated by an MLP hyper-network. We detail the architecture for the selection-decoder further in Appendix B.
The edge choice from the first decoder is fed back into the encoder, where the embeddings of corresponding edges are aggregated into node embeddings using (7). After this final GREAT layer, we find it beneficial to add a small number of standard transformer layers to obtain more expressive node embeddings for the second decoder, which constructs routes.
The routing-decoder is a preference-conditioned version of the multi-pointer decoder (Jin et al. 2023). Similarly to the edge-based version above, given the partial tour , this decoder outputs the probabilities for the next construction step, where now denotes a node and the probability depends on the edge choice .
Remark that when applying GMS-DH to simple graphs, we simply remove the first decoder. Also note that this architecture ensures that the first layers only need to be run once for each instance as they are preference-agnostic and utilize the entire multigraph. This reduces the runtime compared to letting the encoder be preference-aware or pruning the multigraph for each preference before encoding.
From a broader perspective, our approach in GMS-DH can be viewed as a combination of non-autoregressive and autoregressive construction. As a drawback, the edge choice can no longer depend on the context given by the partial tour. Instead, it must be chosen before starting roll-out.
@@ 4.3 Training Algorithm
Our first proposed model, GMS-EB, is trained using the Multi-Objective REINFORCE from Lin, Yang, and Zhang (2022). Given a preference and instance , the model parameterizes a policy according to (9). A tour is associated with a Chebyshev reward corresponding to the negative scalarized cost (5):
The objective during training is to maximize
To stabilize the training, we form a baseline by rolling out tours for the same instance and averaging the reward, according to the POMO framework (Kwon et al. 2020). The policy gradient is then utilized to update the model weights and it is approximated, given one preference , according to
@@
For the second model, the training is more involved. Given an edge selection from the selection-head according to (11), the routing-head parameterizes a policy and obtains a scalar reward by (12) as usual. With a fixed , the objective for this head is to maximize
On the other hand, the task of the selection-head is to ensure that the edge choice is optimal. Let denote the optimal tour on the pruned graph given selection . Then, define the reward and objective function based on the optimal tour according to
@@
We outline the training of GMS-DH in Algorithm 1.
@@
In addition, we utilize a form of curriculum learning to speed up training for both models. We start by training on small graph sizes and then gradually increase the problem size. For GMS-DH in the multigraph setting, we also start by training the routing-head on simple graphs to ensure that the approximation (18) is accurate even during the initial epochs of multigraph training. Further details regarding the curriculum learning can be found in Appendix C.